OpenStudy (anonymous):

amoeba reproduce by splitting in half every 24 hours how many will there be after (a) 8 days and (b) 16 days.

OpenStudy (anonymous):

someone answer my question

OpenStudy (kira_yamato):

N = 2^t (a) N = 2^8 = 256 (b) N = 2^16 = 65536

OpenStudy (alfie):

\[2^8= 256\] \[2^16 = 65 536\]

OpenStudy (anonymous):

how did you guys know to use exponents,

OpenStudy (alfie):

You just have to plug in the number of days in the exponent, "every 24hours" means it splits into two amoebas each day, so if, for example, you would have liked the number of amoebas after four days you would have done: \[2^4\]

OpenStudy (unklerhaukus):

The exponential growth formula \[N(t) = C e^{(t/d)}\] where N(t) is the number at time t C is the initial number and d is the doubling period so in this case C = 1 amoeba d = 1 day